Maria flips a dime four times What is the probability that she will get at least two heads?
It is 0.6875
What is the probability that a coin flipped four consecutive times will always land on heads?
The probability that a coin flipped four consecutive times will always land on heads is 1 in 16.
Since the events are sequentially unrelated, take the probability of heads in 1 try, 0.5, and raise that to the power of 4...
1 in 24 = 1 in 16
Approx 0.0027
What does standard deviation tell about distribution and varity?
It is a measure of the spread of the distribution. The greater the standard deviation the more variety there is in the observations.
P(at least two hits out of 6)=1-P(1 or fewer hits out of 6)=1-binomcdf(6, .3, 1)=1-.420175=0.579825
What is probability of getting an honour card?
In a pack of cards, Ace, King, Queen, and Jack are called honours. There are 16 honours in a pack . An honour can be drawn in 16C1 ways (i.e., n(E)=16).
One card can be drawn from a pack in 52C1 ways.
Required probability is 16/52=4/13.
Can you deal cards to your right in euchre?
No, in euchre, and in almost all card games, you deal to your left.
Independent events.
How many times must you throw two dice to be sure that you get the same sum at least twice?
Since there are 11 different outcomes it is possible that the first eleven throws are all different. But the 12th time you throw must repeat one of the previous results.
If you have 3 fair dice each numbered 1 to 6 what is the chance all three dice show even numbers?
Probability is 1/8 or 0.125 or 12.5%
Explanation:
A fair dice with six sides (numbers) has a probability of 1/2 of rolling an even number. Note: 1/2 = 3/6, where the dice has 3 even numbers out of six total sides.
The probability of three same dice all roll an even number is the probability of one of these (1/2) to the power 3, or (1/2) * (1/2) * (1/2) which equals 1/8.
Why does the formula for standard deviation have a square root in it?
The formula for standard deviation has both a square (which is a power of 2) and a square-root (a power of 1/2).
Both must be there to balance each other, to keep the standard deviation value's magnitude similar to (having the same units as) the sample numbers from which it's calculated. If either is removed from the formula, the resulting standard deviation value will have different units, reducing its usefulness as a meaningful statistic.
What is the Possible result of a probability event called?
It Is Called An Outcome :) ( YOur Welcome)
A pair of dice is rolled what is the probability of getting a sum of 5?
since rolling a dice can give you 36 possible combinations, there are 1/9 chances that you will get a sum of 5.
1+4
2+3
3+2
4+1
the 4 equations above will produce a sum of 5
so,the probability is 4/36, or 1/9.
What is the probability of flipping 25 heads in a row?
The best way to think about this is the following way:
What is the probability of flipping heads once?
1/2
What is the probability of flipping heads twice?
1/4 (1/2 * 1/2)
Using this we can derive the equation to find the probability of flipping heads any number of times. 1/2n
Using this we plug in 25 for n and get
1/225 or as a decimal 2.98023224 x 10-8
or as odds 1:33,554,432
If a coin is tossed 5 times what is the probability of getting 5 heads?
If a fair coin is tossed 5 times, the probability of getting 5 heads is:
P(H,H,H,H,H) = (1/2)5 = 1/32 = 0.03125 = 3.125% ≈ 3.1%
Bag with 10 marbles: 3 orange, 5 black, 2 white.
Rephrasing the question.
If a marble is drawn from the bag, then returned to the bag, and a second marble
is drawn, what is the probability that the first marble turns out white and the second
marble black ?
The probability for a marble to come out white from the bag is:
P(W) = 2/10 = 1/5
The probability for a marble to come out black from the bag is:
P(B) = 5/10 = 1/2
The probability for a marble to come out white, put back in the bag and then take again a marble for a second time and turns out to be black is:
P(B2|W1) = (1/5)∙(1/2) = 1/10 = 0.10 = 10 %
Roles of Probability in different fields of life?
When your driving and you have tow routes you can take, one with red lights that shorter, and one without red lights and is longer. You try and figure out the probability that you will hit the lightes all green and make it sooner, or go the long way and not chance it. Also the chance that you'll be in a meeting and your phone will ring, so do you turn the ringer off or chance it.,,,,,,,,,,,,,,,,Ammar Qasim
You would be more likely to pull out a white marble as there are no red marbles in the bag.
Suppose you roll one die find the probability of not rolling an even number?
Not even is odd. Odd is 1,3,& 5 which is 1/2 of the numbers on the die. Therefore the probability of not rolling an even number is 1/2 or 0.5.
If you roll a fair dice and toss a fair coin what is the probability of getting a 2 and a head?
1/6 x 1/2 = 1/12.
1/6 x 1/2 = 1/12.
1/6 x 1/2 = 1/12.
1/6 x 1/2 = 1/12.
If you toss a coin 3 times what is the probability that it will land on heads 3 times?
The probability of getting heads on three tosses of a coin is 0.125. Each head has a probability of 0.5. Since the events are sequentially unrelated, simply raise 0.5 to the power of the number of tosses (3) and get 0.125, or 1 in 8.